OpenStudy (anonymous):

A rainstorm in Portland, Oregon, has wiped out the electricity in about of the households in the city. A management team in Portland has a big meeting tomorrow, and all members of the team are hard at work in their separate households, preparing their presentations. What is the probability that none of them has lost electricity in his/her household? Assume that their locations are spread out so that loss of electricity is independent among their households.

OpenStudy (nurali):

Using the binomial distribution: \[P(No Loss)=\left(\begin{matrix}6 \\ 0\end{matrix}\right)0.05^0(1-0.05)^6=0.95^6=?\]